How do you solve this equation 2w+2(2w+3)=48?
Answer by checkley77(12844)

2w+2(2w+3)=48
2w+4w+6=48
6w=48-6
6w=42
w=42/6
w=7 ans.
Proof:
2*7+2(2*7+3)=48
14+2(14+3)=48
14+2*17=48
14+34=48
48=48